OpenStudy (anonymous):

Tickets for a show were available at the full price of $22 if purchased at the door but at a discount price of $15 if bought in advance. $6,860 was collected through the sale of 420 tickets. How many tickets sold were full-price tickets and how many were discount tickets? Show how to solve this problem algebraically. Show work for this problem by defining variable(s), writing appropriate equation(s), and showing the solving process. Write a sentence to answer the question.

OpenStudy (anonymous):

let f=full price tickets let d=discounted tickets total of 420 tickets sold so f+d=420 for a total price of $6860 22f+15d=6860 now solve f+d=420 22f+15d=6860 as a system of equations

OpenStudy (anonymous):

Thanks so much! I could not figure this problem out.